# Self-Exploring Language Models: Active Preference Elicitation for Online Alignment

### v1

@@ Abstract
Preference optimization, particularly through Reinforcement Learning from Human Feedback (RLHF), has achieved significant success in aligning Large Language Models (LLMs) to adhere to human intentions. Unlike offline alignment with a fixed dataset, online feedback collection from humans or AI on model generations typically leads to more capable reward models and better-aligned LLMs through an iterative process. However, achieving a globally accurate reward model requires systematic exploration to generate diverse responses that span the vast space of natural language. Random sampling from standard reward-maximizing LLMs alone is insufficient to fulfill this requirement. To address this issue, we propose a bilevel objective optimistically biased towards potentially high-reward responses to actively explore out-of-distribution regions. By solving the inner-level problem with the reparameterized reward function, the resulting algorithm, named Self-Exploring Language Models (SELM), eliminates the need for a separate RM and iteratively updates the LLM with a straightforward objective. Compared to Direct Preference Optimization (DPO), the SELM objective reduces indiscriminate favor of unseen extrapolations and enhances exploration efficiency. Our experimental results demonstrate that when finetuned on Zephyr-7B-SFT and Llama-3-8B-Instruct models, SELM significantly boosts the performance on instruction-following benchmarks such as MT-Bench and AlpacaEval 2.0, as well as various standard academic benchmarks in different settings. Our code and models are available at https://github.com/shenao-zhang/SELM.
@@ 1 Introduction
Large Language Models (LLMs) have recently achieved significant success largely due to their ability to follow instructions with human intent. As the defacto method for aligning LLMs, Reinforcement Learning from Human Feedback (RLHF) works by maximizing the reward function, either a separate model [44; 5; 18] or reparameterized by the LLM policy [49; 48; 4; 70], which is learned from the prompt-response preference data labeled by humans. The key to the success of alignment is the response diversity within the preference data, which prevents reward models (RMs) from getting stuck in local optima, thereby producing more capable language models.
Offline alignment methods [49; 54] attempt to manually construct diverse responses for fixed prompts [11; 24; 72], which, unfortunately, struggles to span the nearly infinite space of natural language. On the other hand, online alignment follows an iterative procedure: sampling responses from the LLM and receiving feedback to form new preference data for RM training [44; 21]. The former step helps explore out-of-distribution (OOD) regions through randomness in sampling. However, in standard online RLHF frameworks, maximizing the expected reward learned from the collected data is the only objective for the LLM, sampling from which often leads to responses clustered around local optima. This passive exploration mechanism can suffer from overfitting and premature convergence, leaving the potentially high-reward regions unexplored.
To address this issue, we propose an active exploration method for online alignment that elicits novel favorable responses. In its simplest form, an optimism term $\alpha\max_{y}r(x,y)$ is added to the reward-fitting objective (e.g., logistic regression on dataset $\mathcal{D}$), denoted as $-\mathcal{L}_{\text{lr}}$, resulting in a bilevel optimization objective for the reward model $r$:
@@
In this paper, we formulate this idea within the context of online direct alignment, where the LLM is iteratively updated without a separate RM. We first introduce two modifications to the bilevel RM objective in 1.1, namely adding KL constraints and using relative maximum reward. Then we derive a simple LLM training objective by applying the closed-form solution of the inner-level problem and reparameterizing the reward with the LLM policy. The resulting iterative algorithm is called Self-Exploring Language Models (SELM). We show that the policy gradient of SELM is biased towards more rewarding areas. Furthermore, by reducing the chance of generating responses that are assigned low implicit rewards, SELM mitigates the indiscriminate favoring of unseen extrapolations found in DPO [49; 48] and enhances exploration efficiency.
@@
In experiments, we implement SELM using Zephyr-7B-SFT [57] and Llama-3-8B-Instruct [38] as base models. By finetuning solely on the UltraFeedback [11] dataset and using the small-sized PairRM [25] for iterative AI feedback, SELM boosts the performance of Zephyr-7B-SFT and Llama-3-8B-Instruct by a large margin on AlpacaEval 2.0 [14] ($+16.24\%$ and $+11.75\%$ LC win rates) and MT-Bench [71] ($+2.31$ and $+0.32$). SELM also demonstrates strong performance on standard academic benchmarks and achieves higher pairwise LC win rates against the iterative DPO baseline. Our code and models are available at https://github.com/shenao-zhang/SELM.
@@ 2 Background
2.1 Large Language Models
A language model $\pi\in\Delta_{\mathcal{Y}}^{\mathcal{X}}$ typically takes the prompt $x\in\mathcal{X}$ as input and outputs the response $y\in\mathcal{Y}$. Here, $\mathcal{X}$ and $\mathcal{Y}$ are finite spaces of prompts and responses, respectively. Given the prompt $x\in\mathcal{X}$, a discrete probability distribution $\pi(\cdot\mid x)\in\Delta_{\mathcal{Y}}$ is generated, where $\Delta_{\mathcal{Y}}$ is the set of discrete distributions over $\mathcal{Y}$. Modern recipes for training LLMs consist of pre-training and post-training procedures, where during pre-training, LLMs learn to predict the next word on a huge and diverse dataset of text sequences in order to understand the underlying patterns and structures of natural language in an unsupervised manner. The post-training procedure aims to align better to end tasks and human preferences with two phases happening in order: Supervised Fine-Tuning (SFT) and human preference alignment. Here, SFT fine-tunes the pre-trained LLM with supervised learning on high-quality data to follow instructions on downstream tasks and obtain a model $\pi^{\text{SFT}}$. In the following of this paper, we focus mainly on preference alignment.
2.2 Reward Modeling and Preference Optimization
@@
Specifically, a point-wise reward $r(x,y):\mathcal{X}\times\mathcal{Y}\rightarrow\mathcal{R}$ represents the Elo score [16] of the response $y$ given the prompt $x$. Then the preference distribution can be expressed by the Bradley-Terry model that distinguishes between the preferred response $y_{w}$ and the dispreferred response $y_{l}$ given prompt $x$, denoted as $y_{w}\succ y_{l}\mid x$, using the logistic function $\sigma$:
@@
where $h$ denotes the human rater and the expectation is over $h$ to account for the randomness of the choices of human raters we ask for their preference. When provided a static dataset of $N$ comparisons $\mathcal{D}=\{x_{i},y_{w,i},y_{l,i}\}_{i=1}^{N}$, the parameterized reward model can be learned by minimizing the following logistic regression loss:
@@
$\mathcal{J}(\pi;\mathcal{D})=\mathbb{E}_{x\sim\mathcal{D},y\sim\pi(\cdot\mid x)}\bigl{[}r(x,y)\bigr{]}-\beta\mathbb{D}_{\text{KL}}(\pi\,||\,\pi_{\text{ref}}),$ (2.3)
where $\beta$ is a hyperparameter and $\mathbb{D}_{\text{KL}}(\pi\,||\,\pi_{\text{ref}}):=\mathbb{E}_{x\sim\mathcal{D}}[\text{KL}(\pi(\cdot\mid x)\,||\,\pi_{\text{ref}}(\cdot\mid x))]$ is the expected Kullback-Leibler (KL) divergence. An ideal $\pi_{\text{ref}}$ is the policy that helps mitigate the distribution shift issue [49; 21] between the true preference distribution and the policy $\pi$ during the off-policy RL training. Since we only have access to the dataset $\mathcal{D}$ sampled from the unavailable true preference distribution, $\pi_{\text{ref}}$ can be obtained by fine-tuning on the preferred responses in $\mathcal{D}$ or simply setting $\pi_{\text{ref}}=\pi^{\text{SFT}}$ and performing RLHF based on the SFT model.
@@
With the motivation to get rid of a separate reward model, which is computationally costly to train, recent works [49; 4; 70; 57; 17] derived the preference loss as a function of the policy by changing of variables. Among them, DPO [49] shows that when the BT model in (2.2) can perfectly fit the preference, the global optimizers of the RLHF objective in (2.3) and the following loss are equivalent:
@@ 3.1 RM-Free Objective for Active Exploration
Formally, for the reward function $r$, the bilevel optimization problem with optimism is formulated as:
@@
Following the implicit reward formulation in DPO, we reparameterize the reward function with $\theta\in\Theta$ as $\widehat{r}_{\theta}(x,y)=\beta(\log\pi_{\theta}(y\mid x)-\log\pi_{\text{ref}}(y\mid x))$, which is the optimal solution of (2.3) and can express all reward classes consistent with the BT model as proved in [49]. With this change of variable, we obtain the RM-free objective for direct preference alignment with optimism:
@@
$\underbrace{-\beta\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}_{t}}\Bigl{[}\sigma\bigl{(}\widehat{r}_{\theta}(x,y_{l})-\widehat{r}_{\theta}(x,y_{w})\bigr{)}\bigl{(}\nabla_{\theta}\log\pi_{\theta}(y_{w}\mid x)-\nabla_{\theta}\log\pi_{\theta}(y_{l}\mid x)\bigr{)}\Bigr{]}}_{\nabla_{\theta}\mathcal{L}_{\text{DPO}}(\pi_{\theta};\mathcal{D}_{t})}$
@@
We note that the second line, corresponding to the gradient of the optimism term, decreases the log-likelihood of response $y$ generated by $\pi_{\theta}$ that has a low value of $\exp(-\widehat{r}_{\theta}(x,y)/\beta)$. Therefore, the added optimism term biases the gradient toward parameter regions that can elicit responses $y$ with high implicit reward $\widehat{r}_{\theta}$, consistent with our intuition outlined in Figure 1.
@@ 3.2 Algorithm
4:     Train the LLM $\pi_{\theta_{t}}=\mathop{\mathrm{argmax}}_{\pi_{\theta}}-\mathcal{L}_{\text{DPO}}(\pi_{\theta};\mathcal{D}_{t})-\alpha\mathbb{E}_{x\sim\mathcal{D}_{t}}[\log\pi_{\theta}(y\mid x)]$ and let $\pi_{\text{ref}}=\pi_{\theta_{t}}$.
@@ 3.3 Self-Exploration Reduces Indiscriminate Favor of Unseen Extrapolations
It has been observed recently [48; 46; 64] that DPO decreases the likelihood of responses generated by the reference policy. It is because for any prompt $x$, at convergence when $\pi_{\theta}\neq\pi_{\text{ref}}$, it holds that
@@
For any $\rho\in\Theta$ in the policy parameter space, let $\widehat{r}_{\rho}(x,y)=\beta(\log\pi_{\rho}(y|x)-\log\pi_{\text{ref}}(y|x))$ be the reparameterized implicit reward. Denote $\pi^{\min}_{\rho}$ as the policy that minimizes the expected implicit reward under the KL constraint, i.e.,
@@
The above theorem states that maximizing the divergence between $\pi_{\theta}$ and $\pi_{\text{ref}}$ is essentially reducing the probability of generating responses with low implicit rewards reparameterized by any policy parameter $\rho$ during training. In other words, the policy not only exploits the existing preference data but also learns to avoid generating the text $y$ that is assigned a low reward value. This process occurs in every iteration with updated reference models. Consequently, responses with high potential rewards are selectively preferred and many commonplace responses receive a small probability mass, thus mitigating the indiscriminate favoring of unseen responses and improving exploration efficiency.
4 Related Work
Data Synthesis for LLMs.
A key challenge for fine-tuning language models to align with users’ intentions lies in the collection of demonstrations, including both the SFT instruction-following expert data and the RLHF preference data. Gathering such data from human labelers is expensive, time-consuming, and sometimes suffers from variant quality [44; 29]. To address this issue, synthetic data [34] has been used for aligning LLMs. One line of work focuses on generating plausible instruction prompts for unlabeled data by regarding the target output as instruction-following responses [31; 60; 27; 55]. Besides, high-quality data can also be distilled from strong models for fine-tuning weaker ones [20; 1; 32; 12; 47]. To construct synthetic datasets for offline RLHF, a popular pipeline [11; 57; 59; 24; 72] involves selecting responses sampled from various LLMs on a set of prompts in the hope to increase the diversity of the data that can span the whole language space. However, data manually collected in such a passive way does not consider what improves the model most through its training, leaving the potentially high-reward regions unexplored.
Iterative Online Preference Optimization
Compared to offline RLHF algorithms [49; 70; 4] that collect preference datasets ahead of training, online RLHF [44; 21], especially the iterative/batched online RLHF [5; 63; 19; 22; 62; 6; 50] has the potential to gather better and better synthetic data as the model improves. As a special case, self-alignment language models align their responses with desired behaviors, such as model-generated feedback [66; 67; 53; 58]. Unfortunately, the above methods still passively explore by relying on the randomness during sampling and easily get stuck at local optima and overfit to the current data due to the vast space of natural language. A notable exception is [15], which proposed to use ensembles of RMs to approximately measure the uncertainty for posterior-sampling active exploration. On the contrary, our method explores based on the optimistic bias and does not estimate the uncertainty explicitly, bypassing the need to fit multiple RMs.
Active Exploration.
In fact, active exploration has been widely studied beyond LLMs. Similar to [15], most existing sample-efficient RL algorithms first estimate the uncertainty of the environment using historical data and then plan with optimism [3; 51; 26], or selecting the optimal action from a statistically plausibly set of action values sampled from its posterior distribution [52; 41; 42; 69]. The proposed self-exploration objective can be categorized as an optimism-based exploration method. However, most previous works require the estimation of the upper confidence bound, which is often intractable. Ensemble methods [43; 8; 37] can serve as approximations to the uncertainty estimation but are still computationally inefficient. MEX [35] proposed to combine estimation and planning in a single objective similar to ours and established theoretical guarantees under traditional RL setups. Concurrent to our work, RPO [36] also proposed to use an adversarially chosen reward model for policy optimization, but focuses on mitigating overoptimization in offline RLHF settings.

### v3

@@ Abstract
Preference optimization, particularly through Reinforcement Learning from Human Feedback (RLHF), has achieved significant success in aligning Large Language Models (LLMs) to adhere to human intentions. Unlike offline alignment with a fixed dataset, online feedback collection from humans or AI on model generations typically leads to more capable reward models and better-aligned LLMs through an iterative process. However, achieving a globally accurate reward model requires systematic exploration to generate diverse responses that span the vast space of natural language. Random sampling from standard reward-maximizing LLMs alone is insufficient to fulfill this requirement. To address this issue, we propose a bilevel objective optimistically biased towards potentially high-reward responses to actively explore out-of-distribution regions. By solving the inner-level problem with the reparameterized reward function, the resulting algorithm, named Self-Exploring Language Models (SELM), eliminates the need for a separate RM and iteratively updates the LLM with a straightforward objective. Compared to Direct Preference Optimization (DPO), the SELM objective reduces indiscriminate favor of unseen extrapolations and enhances exploration efficiency. Our experimental results demonstrate that when fine-tuned on Zephyr-7B-SFT and Llama-3-8B-Instruct models, SELM significantly boosts the performance on instruction-following benchmarks such as MT-Bench and AlpacaEval 2.0, as well as various standard academic benchmarks in different settings. Our code and models are available at https://github.com/shenao-zhang/SELM.
@@ 1 Introduction
Large Language Models (LLMs) have recently achieved significant success largely due to their ability to follow instructions with human intent. As the defacto method for aligning LLMs, Reinforcement Learning from Human Feedback (RLHF) works by maximizing the reward function, either a separate model (Ouyang et al., 2022; Bai et al., 2022; Gao et al., 2023) or reparameterized by the LLM policy (Rafailov et al., 2024b, a; Azar et al., 2023; Zhao et al., 2023), which is learned from the prompt-response preference data labeled by humans. The key to the success of alignment is the response diversity within the preference data, which prevents reward models (RMs) from getting stuck in local optima, thereby producing more capable language models.
Offline alignment methods (Rafailov et al., 2024b; Tang et al., 2024) attempt to manually construct diverse responses for fixed prompts (Cui et al., 2023; Ivison et al., 2023; Zhu et al., 2023), which, unfortunately, struggles to span the nearly infinite space of natural language. On the other hand, online alignment follows an iterative procedure: sampling responses from the LLM and receiving feedback to form new preference data for RM training (Ouyang et al., 2022; Guo et al., 2024). The former step helps explore out-of-distribution (OOD) regions through randomness in sampling. However, in standard online RLHF frameworks, maximizing the expected reward learned from the collected data is the only objective for the LLM, sampling from which often leads to responses clustered around local optima. This passive exploration mechanism can suffer from overfitting and premature convergence, leaving the potentially high-reward regions unexplored.
To address this issue, we propose an active exploration method for online alignment that elicits novel favorable responses. In its simplest form, an optimism term $\alpha\max_{y}r(x,y)$ is added to the reward-fitting objective (e.g., the negative log-likelihood loss $\mathcal{L}_{\text{lr}}$ on dataset $\mathcal{D}$), resulting in a bilevel optimization objective for the reward model $r$:
@@
In this paper, we formulate this idea within the context of online direct alignment, where the LLM is iteratively updated without a separate RM. We first introduce two modifications to the bilevel RM objective in (1.1), namely adding KL constraints and using relative maximum reward. Then we derive a simple LLM training objective by applying the closed-form solution of the inner-level problem and reparameterizing the reward with the LLM policy. The resulting iterative algorithm is called Self-Exploring Language Models (SELM). We show that the policy gradient of SELM is biased towards more rewarding areas. Furthermore, by reducing the chance of generating responses that are assigned low implicit rewards, SELM mitigates the indiscriminate favoring of unseen extrapolations in DPO (Rafailov et al., 2024b, a) and enhances exploration efficiency.
@@
In experiments, we implement SELM using Zephyr-7B-SFT (Tunstall et al., 2023b) and Llama-3-8B-Instruct (Meta, 2024) as base models. By fine-tuning solely on the UltraFeedback (Cui et al., 2023) dataset and using the small-sized PairRM (Jiang et al., 2023) for iterative AI feedback, SELM boosts the performance of Zephyr-7B-SFT and Llama-3-8B-Instruct by a large margin on AlpacaEval 2.0 (Dubois et al., 2024) ($+16.24\%$ and $+11.75\%$ LC win rates) and MT-Bench (Zheng et al., 2024) ($+2.31$ and $+0.32$). SELM also demonstrates strong performance on standard academic benchmarks and achieves higher pairwise LC win rates against the very strong iterative DPO baseline, with almost no additional computational overhead under fair comparisons.
2 Related Work
Data Synthesis for LLMs.
A key challenge for fine-tuning language models to align with users’ intentions lies in the collection of demonstrations, including both the SFT instruction-following expert data and the RLHF preference data. Gathering such data from human labelers is expensive, time-consuming, and sometimes suffers from variant quality (Ouyang et al., 2022; Köpf et al., 2024). To address this issue, synthetic data (Liu et al., 2024a) has been used for aligning LLMs. One line of work focuses on generating plausible instruction prompts for unlabeled data by regarding the target output as instruction-following responses (Li et al., 2023a; Wu et al., 2023; Josifoski et al., 2023; Taori et al., 2023; Li et al., 2024a). Besides, high-quality data can also be distilled from strong models for fine-tuning weaker ones (Gunasekar et al., 2023; Abdin et al., 2024; Li et al., 2023b; Ding et al., 2023; Peng et al., 2023). To construct synthetic datasets for offline RLHF, a popular pipeline (Cui et al., 2023; Tunstall et al., 2023b; Wang et al., 2024b; Ivison et al., 2023; Zhu et al., 2023) involves selecting responses sampled from various LLMs on a set of prompts in the hope to increase the diversity of the data that can span the whole language space. However, data manually collected in such a passive way does not consider what improves the model most through its training, leaving the potentially high-reward regions unexplored.
Iterative Online Preference Optimization.
Compared to offline RLHF algorithms (Rafailov et al., 2024b; Zhao et al., 2023; Azar et al., 2023) that collect preference datasets ahead of training, online RLHF (Ouyang et al., 2022; Guo et al., 2024), especially the iterative/batched online RLHF (Bai et al., 2022; Xu et al., 2023; Chen et al., 2022; Gulcehre et al., 2023; Hoang Tran, 2024; Xiong et al., 2023; Calandriello et al., 2024; Rosset et al., 2024) has the potential to gather better and better synthetic data as the model improves. As a special case, self-aligned models match their responses with desired behaviors, such as model-generated feedback (Yuan et al., 2024; Yuanzhe Pang et al., 2024; Sun et al., 2024; Wang et al., 2024a). Unfortunately, the above methods still passively explore by relying on the randomness during sampling and easily get stuck at local optima and overfit to the current data due to the vast space of natural language. A notable exception is Dwaracherla et al. (2024), which proposed to use ensembles of RMs to approximately measure the uncertainty for posterior-sampling active exploration. On the contrary, our method explores based on the optimistic bias and does not estimate the uncertainty explicitly, bypassing the need to fit multiple RMs.
Active Exploration.
In fact, active exploration has been widely studied beyond LLMs. Similar to Dwaracherla et al. (2024), most existing sample-efficient RL algorithms first estimate the uncertainty of the environment using historical data and then either plan with optimism (Auer, 2002; Russo and Van Roy, 2013; Jin et al., 2020; Mehta et al., 2023; Das et al., 2024), or select the optimal action from a statistically plausibly set of values sampled from the posterior distribution (Strens, 2000; Osband et al., 2013, 2023; Zhang, 2022; Li et al., 2024c). The proposed self-exploration objective can be categorized as an optimism-based exploration method. However, most previous works require the estimation of the upper confidence bound, which is often intractable. Ensemble methods (Osband et al., 2024; Chua et al., 2018; Lu and Van Roy, 2017) can serve as approximations to estimate the uncertainty but are still computationally inefficient.
Concurrent Work.
We highlight the concurrent work (to the first version of the current paper) of Xie et al. (2024); Cen et al. (2024); Liu et al. (2024c), among which Xie et al. (2024) establishes the first analysis of the sample complexity of a DPO algorithm in the online setting of RLHF (formulated as MDPs). All of them focus on incorporating an SFT loss or a similar term (as bonus or penalty) alongside the DPO loss as an optimistic or pessimistic adjustment in the online or offline setting, respectively. Xie et al. (2024); Cen et al. (2024) and the current paper focus on the former, while Liu et al. (2024c) focuses on the latter. In the second version of the current paper, we provide the sample complexity of SELM following the proof technique of Xie et al. (2024). Through a reduction technique from Xie et al. (2024), we show how to connect the sample complexity of SELM to that of existing RL algorithms (Zhong et al., 2022; Liu et al., 2024b), which are not tailored to RLHF but enjoy strong theoretical guarantees.
@@ 3 Background
Large Language Models.
A language model $\pi\in\Delta_{\mathcal{Y}}^{\mathcal{X}}$ typically takes the prompt $x\in\mathcal{X}$ as input and outputs the response $y\in\mathcal{Y}$. Here, $\mathcal{X}$ and $\mathcal{Y}$ are finite spaces of prompts and responses, respectively. Given the prompt $x\in\mathcal{X}$, a discrete probability distribution $\pi(\cdot\mid x)\in\Delta_{\mathcal{Y}}$ is generated, where $\Delta_{\mathcal{Y}}$ is the set of discrete distributions over $\mathcal{Y}$. After pretraining and Supervised Fine-Tuning (SFT), preference alignment is employed to enhance the ability of the language model to follow instructions with human intentions.
@@
Specifically, a point-wise reward $r(x,y):\mathcal{X}\times\mathcal{Y}\rightarrow\mathcal{R}$ represents the Elo score (Elo and Sloan, 1978) of the response $y$ given the prompt $x$. Then the preference distribution can be expressed by the Bradley-Terry model that distinguishes between the preferred response $y_{w}$ and the dispreferred response $y_{l}$ given prompt $x$, denoted as $y_{w}\succ y_{l}\mid x$, using the logistic function $\sigma$:
@@
where $h$ denotes the human rater and the expectation is over $h$ to account for the randomness of the choices of human raters we ask for their preference. When provided a static dataset of $N$ comparisons $\mathcal{D}=\{x_{i},y_{w,i},y_{l,i}\}_{i=1}^{N}$, the parameterized reward model can be learned by minimizing the following negative log-likelihood loss:
@@
$\mathcal{J}(\pi)=\mathbb{E}_{x\sim\mathcal{D},y\sim\pi(\cdot\mid x)}\bigl{[}r(x,y)\bigr{]}-\beta\mathbb{D}_{\text{KL}}(\pi\,||\,\pi_{\text{ref}}),$ (3.3)
where $\beta$ is a hyperparameter and $\mathbb{D}_{\text{KL}}(\pi\,||\,\pi_{\text{ref}}):=\mathbb{E}_{x\sim\mathcal{D}}[\text{KL}(\pi(\cdot\mid x)\,||\,\pi_{\text{ref}}(\cdot\mid x))]$ is the expected Kullback-Leibler (KL) divergence. An ideal $\pi_{\text{ref}}$ is the policy that helps mitigate the distribution shift issue (Rafailov et al., 2024b; Guo et al., 2024) between the true preference distribution and the policy $\pi$ during the off-policy RL training. Since we only have access to the dataset $\mathcal{D}$ sampled from the unavailable true preference distribution, $\pi_{\text{ref}}$ can be obtained by fine-tuning on the preferred responses in $\mathcal{D}$ or simply setting $\pi_{\text{ref}}=\pi^{\text{SFT}}$ and performing RLHF based on the SFT model.
@@
With the motivation to get rid of a separate reward model, which is computationally costly to train, recent works (Rafailov et al., 2024b; Azar et al., 2023; Zhao et al., 2023; Tunstall et al., 2023b; Ethayarajh et al., 2024) derived the preference loss as a function of the policy by changing of variables. Among them, DPO (Rafailov et al., 2024b) shows that when the BT model in (3) can perfectly fit the preference, the global optimizers of the RLHF objective in (3.3) and the following loss are equivalent:
@@ 4.1 RM-Free Objective for Active Exploration
Formally, for the reward $r$, the bilevel optimization problem with optimism is formulated as:
@@
Following the implicit reward formulation in DPO, we reparameterize the reward function with $\theta\in\Theta$ as $\widehat{r}_{\theta}(x,y)=\beta(\log\pi_{\theta}(y\mid x)-\log\pi_{\text{ref}}(y\mid x))$, which is the optimal solution of (3.3) and can express all reward classes consistent with the BT model as proved in (Rafailov et al., 2024b). With the above change of variable, we obtain the RM-free objective for direct preference alignment with optimism:
@@
$\underbrace{\beta\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}_{t}}\Bigl{[}\sigma\bigl{(}\widehat{r}_{\theta}(x,y_{l})-\widehat{r}_{\theta}(x,y_{w})\bigr{)}\bigl{(}\nabla_{\theta}\log\pi_{\theta}(y_{w}\mid x)-\nabla_{\theta}\log\pi_{\theta}(y_{l}\mid x)\bigr{)}\Bigr{]}}_{-\nabla_{\theta}\mathcal{L}_{\text{DPO}}(\pi_{\theta};\mathcal{D}_{t})}$
@@
We note that the second line, corresponding to the gradient of the optimism term, decreases the log-likelihood of response $y$ generated by $\pi_{\theta}$ that has a high value of $\exp(-\widehat{r}_{\theta}(x,y)/\beta)$. Therefore, the added optimism term biases the gradient toward parameter regions that can elicit responses $y$ with high implicit reward $\widehat{r}_{\theta}$, consistent with our intuition outlined in Figure 1.
@@ 4.2 Algorithm
4:     Train the LLM $\pi_{\theta_{t}}=\mathop{\mathrm{argmax}}_{\pi_{\theta}}\{-\mathcal{L}_{\text{DPO}}(\pi_{\theta};\mathcal{D}_{t})-\alpha\mathbb{E}_{x\sim\mathcal{D}_{t}}[\log\pi_{\theta}(y\mid x)]\}$, let $\pi_{\text{ref}}=\pi_{\theta_{t}}$.
@@ 5.1 Self-Exploration Reduces Indiscriminate Favor of Unseen Extrapolations
It has been observed recently (Rafailov et al., 2024a; Pal et al., 2024; Xu et al., 2024) that DPO decreases the likelihood of responses generated by the reference policy. It is because for any prompt $x$, at convergence when $\pi_{\theta}\neq\pi_{\text{ref}}$, it holds that
@@
For any $\rho\in\Theta$ in the policy parameter space, let $\widehat{r}_{\rho}(x,y)=\beta(\log\pi_{\rho}(y\mid x)-\log\pi_{\text{ref}}(y\mid x))$ be the reparameterized implicit reward. Denote $\pi^{\min}_{\rho}$ as the policy that minimizes the expected implicit reward under the KL constraint, i.e.,
@@
The proofs for theorems in this section can be found in Appendix B and C. The above theorem states that maximizing the divergence between $\pi_{\theta}$ and $\pi_{\text{ref}}$ is essentially reducing the probability of generating responses with low implicit rewards reparameterized by any policy parameter $\rho$ during training. In other words, the LLM policy not only exploits the existing preference data but also learns to avoid generating the text $y$ that is assigned a low reward value. This process occurs in every iteration with updated reference models. Consequently, responses with high potential rewards are selectively preferred and many commonplace responses receive a small probability mass, thus mitigating the indiscriminate favoring of unseen responses and improving the exploration efficiency. In the next section, we will formally prove that the self-exploration mechanism is sample-efficient.
5.2 Self-Exploration is Provably Sample-Efficient
Following the proof technique of Xie et al. (2024), we provide the sample efficiency of the self-exploration mechanism by establishing a sublinear cumulative regret. Specifically, the cumulative regret $\mathcal{R}(T)$ up to $T$ iterations is defined as the cumulative performance discrepancy between the learned policy $\pi_{t}$ at iteration $t$ and the optimal policy $\pi^{*}$ over the run of the algorithm:
$\mathcal{R}(T)$ $=\sum_{t=1}^{T}[\mathcal{J}(\pi^{*})-\mathcal{J}(\pi_{t})].$
The key idea is a reduction technique from Xie et al. (2024), which connects the sample complexity of SELM to that of existing RL algorithms (Zhong et al., 2022; Liu et al., 2024b). It is worth noting that the theoretical version of the self-exploration mechanism (Algorithm 2) is a bit different from the practical one used in the numerical experiments and is closer to the proposed algorithm in Xie et al. (2024).
Assumption 5.2 (Realizable Policy Class with Regularity Condition).
We assume access to a policy class $\Pi$ containing the optimal policy $\pi^{*}$. Moreover, we assume that
$\left|\log\frac{\pi(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)}\right|\leq R_{\max}.$
for any $\pi\in\Pi$ and prompt-response pair $(x,y)$.
Assumption 5.2 stipulates that the policy class $\Pi$ is sufficiently comprehensive to include the optimal policy. Additionally, it imposes a bounded condition on $\log(\pi/\pi_{\mathrm{ref}})$, which has been identified as the implicit reward function for DPO (Rafailov et al., 2024b).
Theorem 5.3.
Under Assumption 5.2, let $\eta=\sqrt{Td_{\mathrm{PGEC}}/(\exp(4R_{\max})\log(|\Pi|/\delta))}$, $\alpha=2/(\eta\exp(4R_{\max}))$, and $\delta\in(0,1)$. Then with probability at least $1-\delta$, we have
$\mathcal{R}(T)$ $\lesssim\sqrt{d_{\mathrm{PGEC}}\cdot\exp(2R_{\max})\cdot T\cdot\log(|\Pi|/\delta)},$
where $\lesssim$ omits absolute constants, and $d_{\mathrm{PGEC}}$ is a preference-based version of Generalized Eluder Coefficient (GEC; Zhong et al., 2022) defined in Appendix C.1 capturing the complexity of learning problem. For log-linear policy class $\Pi=\{\pi_{\theta}:\pi_{\theta}(y{\,|\,}x)\propto\exp(\langle\phi(x,y),\theta\rangle/\beta)\}$ with $d$-dimensional feature $\phi$, it holds that $d_{\mathrm{PGEC}}\leq\widetilde{O}(d)$.
The proof technique is from Xie et al. (2024), which connects RLHF with RL and allows us to use the preference-based version of GEC (Zhong et al., 2022; Liu et al., 2024b) as the complexity measure to characterize the cumulative regret $\mathcal{R}(T)$. We restate the proof technique from Xie et al. (2024) for completeness. We emphasize that it is not a novel contribution of the present work. Since the cumulative regret is sublinear in the number of iterations $T$, the above theorem indicates that the policy $\pi_{t}$ converges to the optimal $\pi^{*}$ within sufficient iterations. Moreover, by the standard online-to-batch argument, Theorem 5.3 shows that SELM is capable of finding an $\varepsilon$-optimal policy with a sample complexity of $\widetilde{O}(1/\varepsilon^{2})$. This highlights the sample efficiency of SELM from the theoretical perspective.